Spaces with intrinsically invariant fractal dimensions

Irmina Herburt, Maria Moszyńska and Krzysztof Rudnik, Warsaw, Poland

Keywords: fractal dimension, Hausdorff dimension, box dimension, intrinsic isometry, LGA-space, Lipschitz point AMS Subject classification: 28A80, 54E40, 54E99, 54F45

Abstract: As was proved in our paper [10], in general, a fractal dimension of a metric space does not have to be invariant under intrinsic isometries. The present paper is a continuation of [10]: we study some classes of spaces with intrinsically invariant fractal dimensions.